# Coupling induced emergent topology in a two-leg fermionic ladder

### v1

@@ III.2 Topological Phase Transition
Based on the above analysis of the appearance of zero-energy edge modes under OBC and the quantized topological invariant $\omega$ under PBC, we now characterize the topological properties of our system as a function of $t_{p}$ for different values of dimerization strength (i.e by fixing $t_{2}=1.0$ and varying $t_{1}$).
@@
Figure 5: The phase diagram in the $t_{p}-t_{1}$ plane is shown for $t_{2}=1.0$. In panel (a), the color bar represents the single-particle excitation gap ($G$) at half-filling, computed for a system of size $L=480$ under OBC. In panel (b), the color bar indicates the winding number ($\omega/\pi$) obtained under PBC. The horizontal lines at $t_{1}=t_{2}=1.0$ indicates the gapless line.
@@
To this end, we compute the energy gap ($G$), defined as the energy difference between the two eigenstates at the middle of the bulk gap (i.e., at half-filling), and is given by $G=E_{L}-E_{L-1}$ under OBC. The plot of $G$ (color coded) as a function of $t_{p}$ and $t_{1}$ while fixing $t_{2}=1.0$ is shown in Fig. 5(a). The blue regions, where $G$ vanishes, indicate the presence of degenerate zero-energy modes at half-filling, whereas the in the gray region, $G$ acquires finite values due to the absence of any degenerate edge states. For a fixed $t_{1}$, $G$ remains zero up to a critical value of $t_{p}$, beyond which it becomes finite, suggesting a topological phase transition. The point at which this topological phase transition occurs agrees well with our analytical boundary at $t_{p}=2t(t_{1}+t_{2})$ (indicated by the red line with circles). Note that at $t_{1}=t_{2}$, the entire bulk is gapless up to $t_{p}\leq 2$ indicated by solid cyan line in Fig. 5(a). In Fig. 5(b) we plot the bulk topological invariant $\omega/\pi$ (the winding number) defined in Eq. (14) for the similar parameters as in Fig. 5(a). The values of $\omega/\pi=1~(0)$ confirms the topological (trivial) regions. This clearly complements the topological to trivial phase transition and the gapless line obtained using the gap $G$ in Fig. 5(a). Note that in the gapless region at $t_{1}=t_{2}=1.0$ and up to $t_{p}\leq 2$, the invariant is ill-defined due to the absence of a bulk energy gap and hence the value is not fixed in these parameter regimes. However, the regions of $t_{1}<t_{2}$ and $t_{1}>t_{2}$ are topologically distinct, which we clarify in the following section.
@@
The above analysis demonstrates that a well-defined topological phase emerges in the system due to $t_{p}$, regardless of the dimerization pattern in leg-a, i.e., whether $t_{1}<t_{2}$ or $t_{1}>t_{2}$. Moreover, an increase in the value of $t_{p}$ favours a topological to trivial phase transition for any fixed dimerization strength. In the following, we will quantify the above topological phase transition occurring at half-filling through a quantity that serves as a dynamic probe for characterizing the above topological phase transitions, and also highlight the phenomenon of reversal of Thouless charge pumping.
III.3 Reversal of Thouless charge pumping
@@
Thouless charge pumping (TCP) Thouless (1983) is a versatile platform to characterize the topological phase in the system through periodic modulation of system parameters. This involves slow and cyclic variation of at least two parameters in the system resulting in a quantized particle transport even in the absence of an external bias. One of the parameters drives the system through a topological phase transition, while an additional staggered potential term breaks the inversion symmetry and maintains a finite bulk gap throughout the cycle. This is crucial for ensuring adiabaticity, which is a key requirement for quantized charge transport. During the adiabatic cycle, the system transitions from a topological phase to a trivial phase and back to the initial topological phase, enclosing the topological phase transition point. This process results in the transport of a quantized number of particles and, the number of particles transported is solely determined by the topological invariant, i.e., the Chern number of the system. TCP has been widely investigated in various theoretical and experimental studies in recent years as a key tool for understanding topological phase transitions Lohse et al. (2016); Nakajima et al. (2016); Citro and Aidelsburger (2023); Hayward et al. (2018); Rice and Mele (1982); Citro and Aidelsburger (2023); Asbóth et al. (2016b); Lin et al. (2020); Ke et al. (2017); Argüello-Luengo et al. (2024); Kuno et al. (2017); Bertok et al. (2022); Mondal et al. (2022b); Hayward et al. (2018); Schweizer et al. (2016); Kraus et al. (2012); Taddia et al. (2017); Wang et al. (2013); Walter et al. (2023); Ke et al. (2020); Padhan et al. (2024); Jürgensen et al. (2023, 2021); Viebahn et al. (2024); Zhu et al. (2024). The Rice-Mele model provides a well-established framework for describing charge pumping protocols.
@@
Figure 6: Instantaneous energy spectrum of the Hamiltonian $\hat{H}_{p}$ as a function of the pumping parameter $\tau$, computed for a system of $L=100$ sites under OBC. Panel (a) corresponds to the parameter set $t_{1}=0.2$, $t_{2}=1.0$, while panel (b) shows results for $t_{1}=1.0$, $t_{2}=0.2$. In both cases, the pumping parameters are fixed at $t_{p_{0}}=1.55$, $\delta_{0}=0.5$, and $\Delta_{0}=0.5$. The highlighted markers in (a) and (b) blue circle, green square, and brown star, indicate specific states whose wavefunction amplitudes are shown in panels (c) and (d), respectively. In panel (c), the top, middle, and bottom subplots show the probability density $|\psi_{i}|^{2}$ plotted against the site index for the states marked in (a). Analogously, panel (d) presents the probability density corresponding to the marked states in (b).
@@
To investigate the topological phase transition occurring as a function of $t_{p}$ at half-filling, we propose a pumping protocol that is governed by the Hamiltonian
@@
$\;+\sum_{i}\Delta(\tau)((-1)^{i}\hat{a}_{i}^{\dagger}\hat{a}_{i}+(-1)^{i+1}\hat{b}_{i}^{\dagger}\hat{b}_{i}).$ (8)
@@
To investigate the charge pumping mechanism, we present the instantaneous energy spectrum of the Hamiltonian defined in Eq. (8) for a system of size $L=100$ rungs under OBC. Fig. 6(a) and (b) show the spectra for $t_{1}<t_{2}$ and $t_{1}>t_{2}$, respectively. The cyclic variation of the pumping parameters is represented by cycle-2 (highlighted by a green circle in the insets of Fig. 7(a) and (b)), which encloses the topological phase transition point. In both cases, at half-filling, two in-gap modes cross during the cycle, one ascending from the lower to the upper band and the other descending in the opposite direction after a certain value of the pumping parameter. To demonstrate their edge nature and the evolution during the pump, we plot the probability density of these modes just before and after the half-cycle (i.e. $\tau/2\pi=0.5$), which correspond to the green square and brown star marked points in Figs. 6(a) and (b), respectively. The wave function amplitudes are shown in the middle and lower panels of Figs. 6(c) and (d), respectively. These plots confirm that these modes exchange their position from one edge to another edge across the cycle, realizing topological pumping from one edge site to another edge site of the chain. For $t_{1}<t_{2}$, the edge mode initially localized at the right end of the chain moves to the left after $\tau/2\pi=0.5$, and vice versa for $t_{1}>t_{2}$. This clearly illustrates the reversal of the topological charge pumping direction, governed by the relative dimerization pattern of $t_{1}$ and $t_{2}$. Since the cycle starts from a trivial phase, initially all states near half-filling have finite probability in the bulk, which is shown in the upper panels of Fig. 6(c) and (d) corresponding to the blue circle in Fig. 6(a) and (b), respectively.
Figure 8: (a)The evolution of polarization ($P(\tau)$) as a function of $\tau$ at $t_{p}=2$ cut of the phase diagram shown in Fig. 5 for three different values of $t_{1}$ i.e., $0.5$ (blue squares), $1.0$ (green circles), and $1.5$ (red diamonds). The pumping parameters are set to be $t_{p_{0}}=2.0,\delta_{0}=1.0~~\text{and},~\Delta_{0}=0.6$ (b) The amount of charge pumped ($Q$) by varying $t_{1}$ and fixing the value of $t_{p}$ at $2.0$.
@@
It is evident from Fig. 7(a) that only the cycle 2 results in a smooth variation of $P(\tau)$ from 0 to $1$, corresponding to a quantized charge transfer of $Q=+1$. For the cycle-1, no pumping occurs, while the cycle-3 shows discontinuity in $P(\tau)$, indicating a breakdown of transport. We further explore the scenario by switching the dimerization to $t_{1}=1.0$, $t_{2}=0.2$, while keeping other parameters fixed. In this case, $P(\tau)$ evolves from $1$ to 0 for cycle-2, implying a reversed pumping direction with Q=1Q=-1 [see Fig. 7(b)]. This confirms that the direction of charge transport depends on the dimerization pattern, and altering the dimerization pattern reverses the direction of pumping. To further illustrate the reversal of charge pumping, we fix the rung hopping to $t_{p}=2$ in the phase diagram [Fig. 5], and examine the evolution of polarization $P(\tau)$ for three different values of $t_{1}$: 0.5, 1.0, and 1.5, while keeping $t_{2}=1$ fixed [Fig. 8(a)]. For $t_{1}=0.5$ ($t_{1}<t_{2}$), the polarization smoothly changes from 0 to $1$ over the cycle, indicating a pumped charge of $Q=+1$. At $t_{1}=1.0$, the pumping path intersects the gap-closing point, leading to a breakdown of adiabatic transport and resulting in no net pumping. In contrast, for $t_{1}=1.5$ ($t_{1}>t_{2}$), the polarization evolves from $1$ to 0, corresponding to a pumped charge of Q=1Q=-1, demonstrating a reversal in the direction of charge transport compared to the $t_{1}<t_{2}$ case. To quantify this reversal more generally, we compute the total pumped charge $Q$ as a function of $t_{1}$, keeping all other parameters fixed as in Fig. 8(a). The resulting behavior, shown in Fig. 8(b), clearly shows that the amount of charge pumped changes from $1$ to 1-1 as soon as one crosses the $t_{1}=t_{2}$ point. This behavior captures the transition from positive to negative charge pumping, confirming the reversal of the charge pumping. From the above analysis, it is clear that we find a non-trivial phenomenon of reversal of Thouless charge pumping by periodic modulation of the rung hopping strength. This phenomenon of reversal of TCP also indicates that the phases for $t_{1}<t_{2}$ and $t_{1}>t_{2}$ are topologically distinct. In the following section, we show the many-body properties of the model [Eq. (1)] under consideration.
@@
IV Interaction induced topological phase transition
We now investigate how this topological phase arising in the non-interacting limit evolves upon introducing inter-chain (rung) interactions $V_{p}$. Before analyzing the evolution of topological properties, we first present the bulk phase diagram in the $V_{p}-\mu$ plane, with $t_{1}=0.2,~t_{2}=1.0$ and $t_{p}=0.8$, which is depicted in Fig. 9. The white regions are the gapped phases, and the shaded regions are the gapless phases. The black dashed lines are the edge states, which are extracted by tracking the vanishing up of the single-particle excitation gap defined as $G=\mu_{1}^{+}-\mu_{1}^{-}$, where
@@ IV Interaction induced topological phase transition
Figure 10: The change in density ($\rho$) as a function of change in chemical potential ($\mu$) for different values of $V_{p}$ for a system of $L=100$ rungs. The plateaus at different filling ($\rho=1/4,~1/2,~\text{and}~3/4$) indicate the gap in the system and the length of the plateaus give an estimate of the gaps.
@@
Figure 11: The figure shows the magnitude of bond energies of all the bonds defined in Eq. (13) and the on-site particle number ($\langle\hat{n}_{i}\rangle$) for different phases corresponding to Fig. 9 with a system consisting of $L=100$ rungs. In the figures, the thickness of a bond is proportional to the respective magnitude of the bond energies, and the face color of the circles represents the values of $\langle\hat{n}_{i}\rangle$. (a) The magnitude of bond energy and the onsite particle density at $1/2$ filling ($N=L$) for $V_{p}=0.2$, (b) $V_{p}=2.8$, (c) represents for $1/4$ filling ($N=L/2$) at $V_{p}=1.4$.
To gain further insight into the nature of these gapped phases, we analyze the many-body ground state by examining the magnitude of the NN bond energies and onsite particle densities at different fillings $\rho=1/4,1/2,\text{and}~3/4$, as shown in the phase diagram in Fig. 9. To do this, we compute the bond energies on the legs ($B_{\alpha,i}$) and rungs ($B_{r,i}$) using the following expressions:
$B_{\alpha,i}=\langle\alpha_{i}^{\dagger}\alpha_{i+1}+\text{H.c.}\rangle,\quad B_{r,i}=\langle a_{i}^{\dagger}b_{i}+\text{H.c.}\rangle$ (13)
@@
[…]
@@
Figure 12: On-site particle density for (a) $V_{p}=0.2$ and (b) $V_{p}=2.8$. The face color of the circles represents the onsite particle density. To highlight the presence of edge states, only a few rungs near the boundary of a ladder with $L=100$ rungs are displayed.
Figure 13: (a) shows the winding number $\gamma/\pi$ (squares) calculated using ED for a system of $L=12$ rungs and the string order parameter $O_{\text{str}}$ (circles) computed using DMRG for a system of $L=100$ rungs as a function of $V_{p}$ with the other parameters set to $t_{1}=0.2,~t_{2}=1.0,~\text{and}~t_{p}=0.8$. (b) shows the lowest four values of the entanglement spectrum ($\xi_{1},~\xi_{2},~\xi_{3},~\xi_{4}$) as a function of $V$, for a system of size $L=200$ rungs.
@@
$O_{\text{str}}(r)=-\langle Z_{\text{rung}}(i)e^{i\frac{\pi}{2}\sum_{j=i+1}^{k-1}Z_{\text{rung}}(j)}Z_{\text{rung}}(k)\rangle,$ (15)
where $Z_{\text{rung}}(i)=Z_{a}(i)+Z_{b}(i)$ and $Z_{\alpha}(i)=1-2\alpha_{i}^{\dagger}\alpha_{i}$. Here, $i$ denotes the $i^{th}$ rung of the ladder and r=|ik|r=|i-k| represents the distance between the $i^{th}$ and $k^{th}$ rung. To avoid edge rungs, we have chosen $i=2$ and k=L1k=L-1. We compute $O_{\text{str}}$ for a system of size $L=100$ rungs and plot it along with the winding number in Fig. 13(a) (red circles). A comparison between the behavior of $O_{\text{str}}$ and $\gamma/\pi$, reveals that $O_{\text{str}}$ agrees well with the topological phase transition point identified from the winding number.
@@
V Effect of uniform nearest-neighbor interaction

### v2

@@ III.1.2 Phase diagram
Based on the above analysis of the appearance of zero-energy edge modes under OBC and the quantized topological invariant $\omega$ under PBC, we now demonstrate the topological phase transitions in the form of phase diagrams. For this purpose, we characterize the topological properties of the system as a function of $t_{p}$ for different values of dimerization strength, that is by fixing $t_{2}=1.0$ and varying $t_{1}$.
@@
Figure 5: The phase diagram in the $t_{p}-t_{1}$ plane is shown for $t_{2}=1.0$. In panel (a), the color bar represents the single-particle excitation gap ($G$) at half-filling, computed for a system of size $L=480$ under OBC. In panel (b), the blue regions correspond to a winding number $\omega/\pi=1$ and gray regions to $\omega/\pi=0$ obtained under PBC. The horizontal lines at $t_{1}=t_{2}=1.0$ indicate the gapless line.
@@
To this end, we compute the energy gap ($G$), defined as the energy difference between the two eigenstates at the middle of the bulk gap (i.e., at $\rho=1/2$), and is given by $G=E_{L}-E_{L-1}$ under OBC. The plot of $G$ (color coded) as a function of $t_{p}$ and $t_{1}$ while fixing $t_{2}=1.0$ is shown in Fig. 5(a). The blue regions, where $G$ vanishes, indicate the presence of degenerate zero-energy modes at $\rho=1/2$, whereas in the gray region, $G$ acquires finite values due to the absence of any degenerate edge states. For a fixed $t_{1}$, $G$ remains zero up to a critical value of $t_{p}$, beyond which it becomes finite, suggesting a topological phase transition. The point at which this topological phase transition occurs agrees well with the analytical boundary at $t_{p}=\sqrt{2t(t_{1}+t_{2})}$ (indicated by the red line with circles). Note that at $t_{1}=t_{2}$, the entire bulk is gapless up to $t_{p}\leq 2$ indicated by the horizontal solid line in Fig. 5(a). In Fig. 5(b) we plot the bulk topological invariant $\omega/\pi$ (the winding number) defined in Eq. (14) for the same parameters as in Fig. 5(a). The values of $\omega/\pi=1~(0)$ confirms the topological (trivial) regions. This clearly complements the topological to trivial phase transition as a function of $t_{p}$. Note that in the gapless region at $t_{1}=t_{2}=1.0$ and up to $t_{p}\leq 2$, the invariant is ill-defined due to the absence of a bulk energy gap and hence the value is not fixed in these parameter regimes.
@@
The above analysis demonstrates that a well-defined topological phase emerges in the system due to $t_{p}$, regardless of the dimerization pattern in leg-a, i.e., whether $t_{1}<t_{2}$ or $t_{1}>t_{2}$. Moreover, an increase in the value of $t_{p}$ favours a topological to trivial phase transition for any fixed dimerization strength. In the following, we will quantify the above topological phase transition occurring at $\rho=1/2$ through a quantity that serves as a dynamic probe for characterizing the above topological phase transitions, and also highlight the phenomenon of reversal of Thouless charge pumping.
III.1.3 Reversal of Thouless charge pumping
@@
Thouless charge pumping (TCP) Thouless (1983) is a versatile platform to characterize the topological phase in the system through periodic modulation of system parameters. This involves slow and cyclic variation of at least two parameters in the system resulting in a quantized particle transport even in the absence of an external bias. One of the parameters drives the system through a topological phase transition, while an additional staggered potential term breaks the inversion symmetry and maintains a finite bulk gap throughout the cycle. This is crucial for ensuring adiabaticity, which is a key requirement for quantized charge transport. During the adiabatic cycle, the system traverses from a topological phase to a trivial phase and back to the initial topological phase, enclosing the topological phase transition point. This process results in the transport of a quantized number of particles and, the number of particles transported is solely determined by the topological invariant, i.e., the Chern number of the system. TCP has been widely investigated in various theoretical and experimental studies in recent years as a key tool for understanding topological phase transitions Lohse et al. (2016); Nakajima et al. (2016); Citro and Aidelsburger (2023); Hayward et al. (2018); Rice and Mele (1982); Citro and Aidelsburger (2023); Asbóth et al. (2016b); Lin et al. (2020); Ke et al. (2017); Argüello-Luengo et al. (2024); Kuno et al. (2017); Bertok et al. (2022); Mondal et al. (2022b); Hayward et al. (2018); Schweizer et al. (2016); Kraus et al. (2012); Taddia et al. (2017); Wang et al. (2013); Walter et al. (2023); Ke et al. (2020); Padhan et al. (2024); Jürgensen et al. (2023, 2021); Viebahn et al. (2024); Zhu et al. (2024). The Rice-Mele model provides a well-established framework for describing charge pumping protocols.
@@
Figure 6: Instantaneous energy spectrum of the Hamiltonian $\hat{H}_{p}$ as a function of the pumping parameter $\tau$, computed for a system of $L=50$ rungs under OBC. Panel (a) corresponds to the parameter set $t_{1}=0.2$, $t_{2}=1.0$, while panel (b) shows results for $t_{1}=1.0$, $t_{2}=0.2$. In both cases, the pumping parameters are fixed at $t_{p_{0}}=1.55$, $\delta_{0}=0.5$, and $\Delta_{0}=0.5$. The highlighted markers in (a) and (b) blue circle, green square, and brown star, indicate specific states whose probability amplitudes are shown in panels (c) and (d), respectively. In panel (c), the top, middle, and bottom subplots show the probability density $|\psi_{i}|^{2}$ plotted against the site index for the states marked in (a). Analogously, panel (d) presents the probability density corresponding to the marked states in (b).
@@
To investigate the topological phase transition occurring as a function of $t_{p}$ at $\rho=1/2$, we propose a pumping protocol that is governed by the Hamiltonian
@@
$\;+\sum_{i}\Delta(\tau)((-1)^{i+1}\hat{a}_{i}^{\dagger}\hat{a}_{i}+(-1)^{i}\hat{b}_{i}^{\dagger}\hat{b}_{i}).$ (8)
@@
To investigate the charge pumping mechanism, we present the instantaneous energy spectrum of the Hamiltonian defined in Eq. (8) for a system of size $L=100$ rungs under OBC. Fig. 6(a) and (b) show the spectra for $t_{1}<t_{2}$ and $t_{1}>t_{2}$, respectively. The cyclic variation of the pumping parameters is represented by cycle-2 (highlighted by a green circle in the insets of Fig. 7(a) and (b)), which encloses the topological phase transition point. In both cases, at $\rho=1/2$, two in-gap modes cross during the cycle, one ascending from the lower to the upper band and the other descending in the opposite direction after a certain value of the pumping parameter. To demonstrate their edge nature and the evolution during the pump, we plot the probability density of these modes just before and after the half-cycle (i.e. $\tau/2\pi=0.5$), which correspond to the green square and brown star marked points in Figs. 6(a) and (b), respectively. The probability amplitudes are shown in the middle and lower panels of Figs. 6(c) and (d), respectively. These plots confirm that these modes exchange their position from one edge to another edge across the cycle, realizing topological pumping from one edge site to another edge site of the chain. For $t_{1}<t_{2}$, the edge mode initially localized at the right end of the chain moves to the left after $\tau/2\pi=0.5$, and vice versa for $t_{1}>t_{2}$. This clearly illustrates the reversal of the topological charge pumping direction, governed by the relative dimerization pattern of $t_{1}$ and $t_{2}$. Since the cycle starts from a trivial phase, initially all states near half-filling have finite probability in the bulk, which is shown in the upper panels of Fig. 6(c) and (d) corresponding to the blue circle in Fig. 6(a) and (b), respectively.
Figure 8: (a)The evolution of polarization ($P(\tau)$) as a function of $\tau$ at $t_{p_{0}}=2$ for three different values of $t_{1}$ i.e., $0.5$ (blue squares), $1.0$ (green circles), and $1.5$ (red diamonds), keeping $t_{2}=1$ (Refer Phase diagram shown in Fig. 5). The pumping parameters are set to be $t_{p_{0}}=2.0,\delta_{0}=1.0~~\text{and},~\Delta_{0}=0.6$ (b) The amount of charge pumped ($Q$) by varying $t_{1}$ with fixed value of $t_{p_{0}}=2.0$.
@@
It is evident from Fig. 7(a) that only the cycle 2 results in a smooth variation of $P(\tau)$ from 0 to $1$, corresponding to a quantized charge transfer of $Q=+1$. For the cycle-1, no pumping occurs, while the cycle-3 shows discontinuity in $P(\tau)$, indicating a breakdown of transport. We further explore the scenario by switching the dimerization to $t_{1}=1.0$, $t_{2}=0.2$, while keeping other parameters fixed. In this case, $P(\tau)$ evolves from $1$ to 0 for cycle-2, implying a reversed pumping direction with Q=1Q=-1 [see Fig. 7(b)]. This confirms that the direction of charge transport depends on the dimerization pattern, and altering the dimerization pattern reverses the direction of pumping. To further illustrate the reversal of charge pumping, we fix the rung hopping to $t_{p}=2$ in the phase diagram [Fig. 5], and examine the evolution of polarization $P(\tau)$ for three different values of $t_{1}$: 0.5, 1.0, and 1.5, while keeping $t_{2}=1$ fixed [Fig. 8(a)]. For $t_{1}=0.5$ ($t_{1}<t_{2}$), the polarization smoothly changes from 0 to $1$ over the cycle, indicating a pumped charge of $Q=+1$. At $t_{1}=1.0$, the pumping path intersects the gap-closing point, leading to a breakdown of adiabatic transport and resulting in no net pumping. In contrast, for $t_{1}=1.5$ ($t_{1}>t_{2}$), the polarization evolves from $1$ to 0, corresponding to a pumped charge of Q=1Q=-1, demonstrating a reversal in the direction of charge transport compared to the $t_{1}<t_{2}$ case. To quantify this reversal more generally, we compute the total pumped charge $Q$ as a function of $t_{1}$, keeping all other parameters fixed as in Fig. 8(a). The resulting behavior, shown in Fig. 8(b), clearly shows that the amount of charge pumped changes from $1$ to 1-1 as soon as one crosses the $t_{1}=t_{2}$ point. This behavior captures the transition from positive to negative charge pumping, confirming the reversal of the charge pumping. From the above analysis, it is clear that we find the phenomenon of reversal of Thouless charge pumping by periodic modulation of the rung hopping strength. This phenomenon of reversal of TCP also indicates that the phases for $t_{1}<t_{2}$ and $t_{1}>t_{2}$ are topologically distinct.
Note that such topological phases do not exist for hardcore bosonic ladder, and it is solely due to fermionic statistics.
III.2 Interacting case
We now investigate how this topological phase arising in the non-interacting limit evolves upon introducing interactions in the system. We first explore the effect of inter-chain (rung) repulsive interactions $V_{p}$ and then the effect of repulsive interactions along all the bonds of the system.
@@ III.2 Interacting case
III.2.1 Effect of inter-chain interaction
Before analyzing the evolution of topological properties, we first present the bulk phase diagram in the $V_{p}-\mu$ plane, with $t_{1}=0.2,~t_{2}=1.0$ and $t_{p}=0.8$, which is depicted in Fig. 9. The white regions are the gapped phases, and the shaded regions are the gapless phases. The black dashed lines are the edge states, which are extracted by tracking the vanishing up of the single-particle excitation gap defined as $G=\mu_{1}^{+}-\mu_{1}^{-}$, where
@@ III.2.1 Effect of inter-chain interaction
To gain further insight into the nature of these gapped phases, we analyze the many-body ground state by examining the magnitude of the NN bond energies and onsite particle densities at different fillings $\rho=1/4,1/2,~\text{and}~3/4$, as shown in the phase diagram in Fig. 9. To do this, we compute the bond energies on the legs ($B_{i,\alpha}$) and rungs ($B_{i,r}$) using the following expressions:
$B_{i,\alpha}=\langle\hat{\alpha}_{i}^{\dagger}\hat{\alpha}_{i+1}+\text{H.c.}\rangle,\quad B_{i,r}=\langle\hat{a}_{i}^{\dagger}\hat{b}_{i}+\text{H.c.}\rangle$ (13)
@@
Figure 10: The change in particle filling ($\rho$) as a function of change in chemical potential ($\mu$) for different values of $V_{p}$ for a system of $L=100$ rungs keeping $t_{1}=0.2,~t_{2}=1.0,~\text{and}~t_{p}=0.8$ fixed. The plateaus at different filling ($\rho=1/4,~1/2,~\text{and}~3/4$) indicate the gap in the system and the length of the plateaus give an estimate of the gaps.
Building on this understanding of the bulk phases, we now examine how the topological characteristics identified in the non-interacting limit evolve as a function of rung interaction. Note that the black dashed lines in Fig. 9 are the edge states, which also existed in the non-interacting limit (see Fig. 3). We obtain that these edge states also survive in the presence of interaction and at $\rho=1/2$, with an increase in $V_{p}$, a gap-closing phase transition occurs to a trivial phase at $V_{p}\sim 2.3$, which we call the interaction-induced topological phase transition. To show that these states exist at the edges in the presence of finite interaction as well, we plot the on-site particle density of both the legs as a function of the rung index for $V_{p}=0.2$ in Fig. 12(a). For comparison, we also show the onsite particle density $\langle\hat{n}_{i,\alpha}\ranglecorresponding$ to the trivial phase, i.e., at $V_{p}=2.8$, which does not have any edge states as shown in Fig. 12(b). The clear asymmetry in density observed in both the edge sites for $V_{p}=0.2$ confirms the edge states. In contrast, for $V_{p}=2.8$, the density remains uniform throughout, confirming that there are no edge states in the system.
@@
Figure 11: The figure shows the magnitude of bond energies of all the bonds defined in Eq. (13) and the on-site particle density for different phases corresponding to Fig. 9 with a system consisting of $L=100$ rungs keeping $t_{1}=0.2,~t_{2}=1.0,~\text{and}~t_{p}=0.8$ fixed. In the figures, the thickness of a bond is proportional to the respective magnitude of the bond energies, and the face color of the circles represents the values of $\langle\hat{n}_{i,\alpha}\rangle$. (a) The magnitude of bond energy and the onsite particle density at $1/2$ filling ($N=L$) for $V_{p}=0.2$, (b) $V_{p}=2.8$, and (c) represents for $1/4$ filling ($N=L/2$) at $V_{p}=1.4$.
Figure 12: On-site particle density for (a) $V_{p}=0.2$ and (b) $V_{p}=2.8$ keeping $t_{1}=0.2,~t_{2}=1.0,~\text{and}~t_{p}=0.8$ fixed. The face color of the circles represents the onsite particle density $\langle\hat{n}_{i,\alpha}\rangle$. To highlight the presence of edge states, only a few rungs near the boundary of a ladder with $L=100$ rungs are displayed.
Figure 13: (a) shows the winding number $\gamma/\pi$ (squares) calculated using ED for a system of $L=12$ rungs and the string order parameter $O_{\text{str}}$ (circles) computed using DMRG for a system of $L=100$ rungs as a function of $V_{p}$ with the other parameters set to $t_{1}=0.2,~t_{2}=1.0,~\text{and}~t_{p}=0.8$. (b) shows the lowest four values of the entanglement spectrum ($\xi_{1},~\xi_{2},~\xi_{3},~\xi_{4}$) as a function of $V_{p}$, for a system of size $L=200$ rungs.
@@
$O_{\text{str}}(r)=-\langle\hat{Z}_{\text{rung}}(i)e^{i\frac{\pi}{2}\sum_{j=i+1}^{k-1}\hat{Z}_{\text{rung}}(j)}\hat{Z}_{\text{rung}}(k)\rangle,$ (15)
where $\hat{Z}_{\text{rung}}(i)=\hat{Z}_{a}(i)+\hat{Z}_{b}(i)$ and $\hat{Z}_{\alpha}(i)=1-2\hat{\alpha}_{i}^{\dagger}\hat{\alpha}_{i}$. Here, $i$ denotes the $i^{th}$ rung of the ladder and r=|ik|r=|i-k| represents the distance between the $i^{th}$ and $k^{th}$ rung. To avoid edge rungs, we have chosen $i=2$ and k=L1k=L-1. We compute $O_{\text{str}}$ for a system of size $L=100$ rungs and plot it along with the winding number in Fig. 13(a) (red circles). A comparison between the behavior of $O_{\text{str}}$ and $\gamma/\pi$, reveals that $O_{\text{str}}$ agrees well with the topological phase transition point identified from the winding number.
@@
III.2.2 Effect of uniform interaction